OpenStudy (anonymous):

IQ scores are normally distributed with a mean of 100 and a standard deviation of 15. If a certain statistician has an IQ of 122, what percent of the population has an IQ less than she does? A. 7% B. 22% C. 93% D. 99% E. 47%

OpenStudy (anonymous):

\(z=\frac{x-\mu}{\sigma}=\frac{122-100}{15}=1.4667\) Then use a z-table to find the percentage.

Directrix (directrix):

C. 93% ---------- Following from above, a Z-score of 1.4667 has area approximately .4279 between the z-score and the mean. To the left of the mean is .5000 area. Added, the area is .9279 which is roughly 93% of the population with IQ less than that of the statistician.

OpenStudy (hba):

ur done with it